OpenStudy (anonymous):

(b) A ball of mass 400 g is thrown with an initial velocity of 30.0 m s–1 at an angle of 45.0° to the horizontal. Air resistance is negligible. The ball reaches a maximum height H after a time of 2.16 s. Calculate 2. the maximum height H of the ball,

OpenStudy (isaiah.feynman):

Use this\[y = u \sin (\theta)t - \frac{ 1 }{ 2 }g t^{2}\]

OpenStudy (anonymous):

other way to solve is , S=0+0.5gt*2, from where zero comes.

OpenStudy (isaiah.feynman):

Nope. You can't have zero, because initial velocity is not zero.

OpenStudy (anonymous):

this answer is given in mark scheme, they gave two different ways to solve

OpenStudy (isaiah.feynman):

What's the other way?

OpenStudy (anonymous):

s=0+0.5(9.81)(2.16)*2

OpenStudy (anonymous):

answer is either 22.88 or 22.94

OpenStudy (isaiah.feynman):

I'd have to say its incorrect.

OpenStudy (anonymous):

ok, thanks for your answer .

OpenStudy (isaiah.feynman):

Nope.

OpenStudy (shamim):

@ali_moazzam is also correct

OpenStudy (shamim):

He consider the object is falling frm H height to ground

OpenStudy (shamim):

So its acceleration

OpenStudy (shamim):

U know Time needed for an object to go frm bottom to top is just equal to time needed to go frm top to bottom

OpenStudy (shamim):

So if we consider the object is going frm top to bottom then the equation will b y=ut*.5gt^2 y=0*t+.5gt^2

OpenStudy (shamim):

For a falling object initial velocity u=0

OpenStudy (shamim):

Anyway answer should b equal. If not then tell me plz!!

OpenStudy (shamim):

I hv 1 more another way to solve this problem. If u wanna c then response plz!!

OpenStudy (anonymous):

no, thanks i understood , thanks !

OpenStudy (isaiah.feynman):

We're wrong. I read wrong. The maximum height is given as \[H = \frac{ u^{2}(\sin \theta)^{2} }{ 2g }\]